# Dual Diffusion for Unified Image Generation and Understanding

### v1

@@ 5.1 Diffusion Models
Diffusion models [67, 29, 68] generate data by gradually converting noise into signal via a reverse diffusion process. They are the de facto standard for image generation [14, 32] and likelihood modeling [33, 55, 69, 38]. Conditional diffusion models [28] have also been shown to be powerful interfaces bridging text and images, particularly for their ability to generate highly realistic and aesthetic images from textual descriptions [63, 64, 57, 22, 10, 62]. Their exceptional performance in the image domain has also inspired numerous extensions to the language generation [41, 12, 15, 52, 27, 51, 65, 23, 66], and is an attractive alternative as its sampling is not constrained by a specified token generation order and the attention mechanism does not need to be uni-directional.
@@ 5.2 Vision Language Models
The success of large language models (LLMs) [9, 76] and vision-language pretraining [60] has given rise to a series of multi-modal language models. The visual signal is projected to the text embedding space via vision encoders supervised by text labels [60, 84] and then connected to a pretrained language models through further instruction tuning [47, 86, 13, 2, 75]. While these models have shown promising capabilities in image understanding and few-shot generalization, their predictive targets are inherently language-centric, limiting their ability to model the image distributions directly.
@@ 5.3 Multimodal Text and Image Generative Models
Limitations
While discrete diffusion offers the advantage of being agnostic to sequential order and is compatible with bi-directional Transformers, its current implementation requires the sequence length to be preset before sampling. A promising future direction would be to extend the sampling scheme to allow for more flexibility, enabling dynamic sequence lengths during the sampling process. In addition, while we show that our proposed dual diffusion model can perform instruction tuning, its instruction-following capabilities still marginally lag behind those of state-of-the-art autoregressive models.
@@ 6 Training Details
Text loss weight 0.2 0.3
@@ 7 Further Results
We evaluate the aesthetic quality of generated images from our proposed D-DiT against those of the original SD3 model and a selection of existing text-to-image (T2I) and multi-modal works. We measure Frechet Inception Distance (FID) with respect to a collection highly aesthetic generated images, known as the MJHQ-30K benchmark proposed by [37]. As shown in Table 7, we observe an improvement in FID after joint diffusion training, and favorable comparison against multi-modal models of similar size.
@@
We also include a qualitative comparison in captioning performance compared to UniDiffuser [7], another diffusion-based multi-modal model, in Figure 6, where we demonstrate an improvement in the ability to capture fine-grained details of the image in a longer caption format. Finally, we provide further uncurated text-to-image (T2I) generation results in Figures 7, 8, 9, and 10. Overall, these results further demonstrate the multi-faceted performance of our proposed dual-branch diffusion-based multi-modal model.
@@
Figure 6: Comparison of captions generated by D-DiT and UniDiffuser[7]. The prompt to D-DiT is ”Provide a brief description of the given image.”

### v2

@@ 5.1 Diffusion Models
Diffusion models [70, 30, 71] generate data by gradually converting noise into signal via a reverse diffusion process. They are the de facto standard for image generation [15, 34] and likelihood modeling [35, 58, 72, 40]. Conditional diffusion models [29] have also been shown to be powerful interfaces bridging text and images, particularly for their ability to generate highly realistic and aesthetic images from textual descriptions [66, 67, 60, 23, 10, 65]. Their exceptional performance in the image domain has also inspired numerous extensions to language generation [43, 12, 16, 55, 28, 54, 68, 24, 69], and is an attractive alternative as its sampling is not constrained by a specified token generation order and the attention mechanism does not need to be uni-directional.
@@ 5.2 Vision Language Models
The success of large language models (LLMs) [9, 79] and vision-language pretraining [63] has given rise to a series of multi-modal language models. The visual signal is projected to the text embedding space via vision encoders supervised by text labels [63, 87] and then connected to a pretrained language model through further instruction tuning [50, 89, 14, 2, 78]. While these models have shown promising capabilities in image understanding and few-shot generalization, their predictive targets are inherently language-centric, limiting their ability to model the image distributions directly.
@@ 6 Training Details
Text loss weight 0.2 1.0
@@ 7 Further Results
Model COCO-30k T2I CompBench
FID $\downarrow$ CLIP $\uparrow$ Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$
SD3 10.2 30.9 0.7993 0.5816 0.7389
D-DiT 9.4 31.2 0.8001 0.5703 0.6856
Table 8: Further image generation comparisons against original SD3 on MS-COCO dataset [47] and T2I CompBench [31].
Image generation
We evaluate the aesthetic quality of generated images from our proposed D-DiT against those of the original SD3 model and a selection of existing text-to-image (T2I) and multi-modal works. We measure Frechet Inception Distance (FID) with respect to a collection highly aesthetic generated images, known as the MJHQ-30K benchmark proposed by [39]. As shown in Table 7, we observe an improvement in FID after joint diffusion training, and favorable comparison against multi-modal models of similar size. We also provide further comparisons on MS-COCO 30k and T2I CompBench in Table 8. The FID and CLIP score slightly improve compared to the original SD3 model. On T2I CompBench, we find that after dual diffusion fine tuning the model performs worse in texture. We hypothesize that the major reason is the texture quality of our training dataset is worse than the dataset used for training SD3.
Text generation process
@@
Model # trainable Text encoder Geneval COCO FID VQAv2(val)
0-shot 32-shot
End-to-End 1.1B - 0.39 18.1 54.3 58.7
From SD3 2B T5-XXL 0.65 9.4 55.0 60.3
Table 9: Comparison of different D-DiT variants. End-to-End variant is trained from scratch and uses GPT2’s text tokenizer. From SD3 variant is initialized from SD3 pretrained checkpoint and uses T5 encoder. The end-to-end model is first trained on OpenWebText for 350B tokens, then trained on DataComp-recap1B for an epoch (400k steps) and a filtered subset for 100k steps.
Training from scratch and removing T5 encoder
To study the influence of text-to-image pretraining, we conduct a study by comparing a D-DiT model that is trained from scratch. We found that initializing from pretrained text-to-image model and use a pretrained text encoder can greatly aid model learning of text-to-image tasks. Meanwhile, image captioning on VQA also mildly improves (Table 9).
Image generation’s influence on SFT
To analyze the influence of dual diffusion loss on image understanding, we conduct supervised finetune on LLaVA 1.5 dataset with varying amount of image generation data, including a training that only has understanding loss (no generation data). We observe that the image generation loss and corresponding data amount does not have significant influence on model’s understanding performance (Table 10).
Und. Gen. VQAv2 (val) POPE
10k 30k 50k 10k 30k 50k
0.665M 0 52.8 55.9 58.3 79.6 80.9 81.8
0.665M 7M 53.4 55.8 58.1 79.8 81.2 82.4
0.665M 20M 53.6 55.8 58.3 81.0 81.1 82.5
Table 10: Understanding performance (accuracy) under different data settings and training steps during supervised finetune. Batch size is set to 128 for this experiment.
Comaprison against previous multi-modal diffusion model
We also include a qualitative comparison in captioning performance compared to UniDiffuser [7], another diffusion-based multi-modal model, in Figure 7, where we demonstrate an improvement in the ability to capture fine-grained details of the image in a longer caption format. Finally, we provide further uncurated text-to-image (T2I) generation results in Figures 8, 9, 10, and 11. Overall, these results further demonstrate the multi-faceted performance of our proposed dual-branch diffusion-based multi-modal model.
(a) T2I Prompt: Three trucks parking in parallel: one red, one blue, and one white. Red truck has load and the rest don’t have.
(b) I2T Prompt:Q: How many pens are there on the desk and what are their colors? A: There are three pens on the desk, and they are red and blue.
Figure 5: Examples of failed text-to-image and image-to-text generation.
Limitations
As shown in Figure 5(b): in T2I, we find that D-DiT can struggle to generate scenes with relatively complex instructions. In I2T, D-DiT can fail to identify the full details of smaller objects. We also observe model’s performance performance deteriorates with longer prompts, primarily due to the bias towards short prompts in the LLaVA finetuning dataset.
In summary, while discrete diffusion offers the advantage of being agnostic to sequential order and is compatible with bi-directional Transformers, its current implementation requires the sequence length to be preset before sampling. A promising future direction would be to extend the sampling scheme to allow for more flexibility, enabling dynamic sequence lengths during the sampling process. In addition, while we show that our proposed dual diffusion model can perform instruction tuning, its instruction-following capabilities still marginally lag behind those of state-of-the-art autoregressive models.
@@
Figure 7: Comparison of captions generated by D-DiT and UniDiffuser[7]. The prompt to D-DiT is "Provide a brief description of the given image."